OpenStudy (anonymous):

State Fair ticket prices for kids under 12 years of age increased from $3.00 to $5.00 this year. What is the approximate percent of increase in the tickets?

OpenStudy (anonymous):

(new - old)/old (5-3)/3 = 2/3 =67% increase